# Model Merging in LLMs, MLLMs, and Beyond: Methods, Theories, Applications and Opportunities

### v4

@@ 4.3.2 Model Merging for Better Domain Generalization
SWAD [17] demonstrates that flatter minima generalize better to unseen domains. Inspired by SWA [73], SWAD proposes a dense and overfit-aware stochastic weight sampling strategy to identify these flatter minima. More specifically, unlike SWA, it starts from a predefined epoch until the final epoch, and collects a random weight every $K$ epochs for averaging. SWAD collects weights densely, that is, one is collected every iteration/step, and the start and end of random weight collection are determined by the performance changes on the validation set. EoA [8] also shows that model averaging can improve out-of-domain performance stability, and that ensembling multiple moving average models can further enhance performance compared to ensembling models without weight averaging.
4.4 Model Merging in Federated Learning
@@
Federated Learning (FL) is a distributed learning approach that allows multiple clients to collaboratively train a model without sharing data. FL primarily includes two settings: centralized (with a central server) and decentralized (without a central server). Each client updates the model or calculates the gradient based on local data and sends the updated information to the central server (in centralized FL) or other clients (in decentralized FL) for aggregation to update the global model, thus ensuring data privacy protection.
4.4.1 Federated Learning Paradigm
Model merging is a routine and crucial operation in FL. Taking centralized FL as an example, it typically involves $N$ clients and a central server $S$. Each client has a private set of training data. Specifically, the training process in the centralized FL paradigm consists of five steps: (1) Model initialization: The central server initializes the global model parameter; (2) Model distribution: The latest model on the server is sent to the local client in the $t$-th round of communication. (3) Update of the local model: The $i$-th client updates the model by calculating the gradient based on the local data. (4) Model upload: The updated models of all local clients are sent to the server aggregation. (5) Model aggregation: The multiple local models on the server are aggregated. These five steps are repeated until the model converges or the maximum number of training rounds is reached. Since this paper is not a survey of FL, we focus on implementing the ‘model aggregation’ step. In FL, model merging refers to summarizing model parameters from various clients during each communication round, thereby forming an updated global model.
4.4.2 Model Merging for Local Knowledge Aggregation
@@
Most FL methods adopt a simple coordinate-wise average to aggregate the local models. For example, they calculate local model merging coefficients according to some heuristic rules. FedAvg [127], the most classic FL method, proposes to merge local models on the server weighted by the amount of training data from each client. FedNova [198] normalizes and scales model updates on the client side based on the number of update steps, efficiently aggregating local models to obtain a high-performance global model. FedAtt [82] calculates layer-wise attention coefficients based on the similarity of client and server parameters, fusing local models based on these coefficients. FedFisher [81] computes the Fisher information matrix of the parameters in each client to merge the local models. In more challenging FL tasks, the above direct coordinate-wise merging methods may result in suboptimal global model performance. Inspired by the property of permutation invariance of neural networks, PFNM [229], OTFusion [167] and FedMA [197] propose to permute neurons of local models before merging them. Similarly, GAMF [112] transforms the model merging problem into a multi-graph matching problem based on graph matching and then merges the aligned local models.
4.5 Model Merging in Zero-shot/Few-shot Learning
@@
In practical applications of machine learning models, collecting a large amount of labeled data can be expensive or infeasible in specific scenarios (e.g., medical diagnosis, real-time monitoring). Users often want deep models to effectively perform new tasks that have not been encountered before, that is, an ability commonly referred to as cross-task generalization [67]. Zero-shot [128] and few-shot learning [224] can reduce the dependence on large amounts of data and allow the model to better deal with unseen categories or small numbers of samples, improving the cross-task generalization ability of the model. In few-shot learning, a common approach is to fine-tune the model using the limited examples available. However, because of the minimal data, this fine-tuning process is often unstable and yields only modest performance improvements. Recently, some studies have explored merging pre-trained models (from some publicly accessible resources) to enhance cross-task generalization under zero-shot and few-shot conditions [211, 67, 79, 137, 129, 196, 181].
@@
4.5.1 Model Merging for Cross-task Generalization in Zero-shot Learning
Model merging has demonstrated the effectiveness of zero-shot learning across several applications. Some examples of practical applications include cross-lingual transfer [69, 27, 236, 95], hybrid style image generation [14, 132], and multi-modal processing [18].
@@ 4.5.1 Model Merging for Cross-task Generalization in Zero-shot Learning
Some works achieve cross-lingual transfer through model merging, such as chat [69], text summarization [27], or reasoning [236]. A well-performing language-specific LLM needs to be fully trained, and with 7,000 languages in the world, not all of them have enough labeled data to support model fine-tuning. Therefore, cross-lingual knowledge transfer is particularly important. For example, Huang et al. [69] build a Chat vector based on fine-tuned LLAMA2-chat and pre-trained LLAMA2 on chat data in the English language, and assembles it with the continuously pre-trained LLAMA2 model on other non-English languages. This allows the new model to chat in non-English languages. Chronopoulou et al. [27] develop a zero-shot multilingual summarization framework. It uses a merged model (a supervised summarization model and an unsupervised pre-trained model for a high-resource language, along with an unsupervised pre-trained model for a low-resource language) to perform text summarization tasks in low-resource languages. Similarly, AdaMergeX [236] demonstrates the effectiveness of model merging for cross-language transfer across three tasks: reasoning, natural language understanding, and natural language generation. In the hybrid style image generation task, Diffusion Soup [14] and MaxFusion [132] show that the zero-shot generation ability can be enhanced by merging multiple diffusion models. In the multi-modality task, DAMC [18] experiments prove that zero-shot multi-modal extension can be achieved by merging multi-modal models, provided they are initialized from the same LLM. For example, by merging a visual LLM and an audio LLM, the combined model can not only perform image or audio tasks independently but also acquire the zero-shot ability to process inputs containing both visual and auditory information simultaneously.
4.5.2 Model Merging for Cross-task Generalization in Few-shot Learning
Parameter-efficient fine-tuning (PEFT), such as LoRA or Adapter, facilitates the creation and sharing of thousands of custom PEFT modules, each trained on different data for various downstream tasks. A natural question is whether combining PEFT modules pre-trained on different upstream tasks can improve the transfer accuracy for unseen downstream tasks with limited samples.
@@ 4.5.2 Model Merging for Cross-task Generalization in Few-shot Learning
Recent work on model merging suggests a positive answer, showing that merged models can enhance generalization in few-shot settings [58, 9, 67, 196]. For example, LoraHub [67] proposes to merge LoRA modules available on HuggingFace to achieve adaptive performance for unseen tasks, where the merging coefficients of different LoRA are searched in a black-box gradient-free manner with few-shot samples. As expected, few-shot LoraHub performs better than few-shot in-context learning and reduces inference costs by eliminating the need for examples as input to LLMs. LoraRetriever [237] further proposes dynamically retrieving the most relevant LoRAs based on the input and merging them. Similarly, MerA [58] proposes merging pretrained adapters into a single adapter for few-shot NLP scenarios. In general, well-trained LoRAs or adapters can serve as valuable resources that users can easily share, access, and apply to a variety of downstream tasks. In the real world, upstream and downstream tasks can be entirely disparate, originating from different datasets, domains, or even different parts of the same dataset. Asadi et al. [9] comprehensively evaluates model merging in the few-shot learning setting. Specifically, this study examines three cases of label, domain, and task drift between upstream and downstream tasks. The results demonstrate that model merging enhances the model’s generalization ability in few-shot learning scenarios across different contexts.
4.6 Model Merging in Adversarial Learning
@@
In the machine learning community, the open-source availability of pre-trained models [189, 188, 184, 145, 146] has accelerated technological advancements. In this context, developers often download unvalidated checkpoints to fine-tune their models or even outsource the training process to third-party platforms [210]. Consequently, open-source models are also vulnerable to malicious attacks, such as poisoning attacks, where hidden malicious behaviors can be triggered by specific inputs. This raises several intriguing questions: Can model merging lead to attacks, and can it be used to develop defense mechanisms? Additionally, how can intellectual property protection be enhanced in the context of model merging?
4.6.1 Model Merging as an Attack Strategy
@@
Parameter-Efficient Fine-Tuning (PEFT) methods [38], such as LoRA [65], exhibit functional transferability. This means that a LoRA model fine-tuned for a specific task based on a pretrained model can be successfully transferred to another pretrained model [116]. In practice, developers often download LoRA models from open-source platforms to address their specific downstream tasks [67]. If a poisoned LoRA, which could be seen as a Trojan horse, is inadvertently downloaded and integrated into a model, it may introduce security vulnerabilities. Research by LoRA-as-an-Attack [116] demonstrates that merging a poisoned LoRA—trained on compromised data—with a benign LoRA, trained on clean data, can result in a backdoor injection. This finding also holds when multiple LoRAs are merged. In addition, BadMerging [235] has developed a two-stage backdoor attack framework specifically for model merging, and through a large number of experiments, it has shown that the success rate of on-task and off-task attacks on merged models exceeds 90%, and existing defense measures cannot defend against BadMerging.
4.6.2 Model Merging as a Defense Strategy
@@
Contrary to the attacks described in §4.6.1, the transferability of LoRA also offers an opportunity for model merging as a defense strategy. Specifically, if we know that a model may be susceptible to certain attacks, can we train some LoRAs to enhance the model’s defense (i.e., reduce the attacker’s success rate)? For example, Liu et al. [116] demonstrate that GPT-3.5 was used to generate a benign dataset containing backdoor triggers. A dedicated defense LoRA was then trained on this benign data and merged into the poisoned pre-trained model. This defensive model merging ultimately led to a reduction in the backdoor effect. Furthermore, research has shown that in the context of full parameter fine-tuning, model merging can serve as a ”free lunch” for model defense. Experiments involving four model architectures and four datasets revealed that merging multiple poisoned models without additional effort mitigated these poisoning attacks, with the accuracy on the benign dataset remaining nearly unaffected. Rebuffi et al. [154] and Croce et al. [30] merge a set of $l_{p}$ (for various $p$) robust fine-tuned models to easily control the robustness level of each threat model against $l_{p}$ boundary adversarial attacks. Similarly, the experimental analysis by [49] indicates that model merging offers an effective defense mechanism against jailbreak attacks [204].
@@
In another practical scenario, merging unauthorized models may infringe on the intellectual property rights of the model owner. Malicious users might merge several high-quality open-source models (e.g., those authorized for research use only) to create a new model, then claim that this new model was entirely developed and trained from scratch by themselves, subsequently offering model services for commercial gain. In such cases, it becomes particularly crucial for model owners to detect whether others have merged their models. MergeGuard [28] performs a preliminary analysis of the effectiveness of two existing defense methods—Quantization Watermarking [102] and Instructional Fingerprint [214]—in the context of model merging. The study observed that while the watermarking method cannot be detected in the merged model, the fingerprint method remains detectable.
5 Remaining Challenges and Future Directions
@@
Although §2, §3 and §4 present various advanced model merging methods and applications, challenges remain in the technology and application of existing model merging approaches. Additionally, there are numerous areas that warrant further research in the future.
@@
(1) Closing the Performance Gap Between the Merged and Independent Models. In practical settings, guaranteeing the performance of model merging remains challenging. The effectiveness of current model merging techniques heavily relies on the ”pretraining-finetuning” paradigm. Specifically, successful model merging requires that multiple models be fine-tuned based on the same pre-trained model, with careful control over the number of epochs and learning rate during fine-tuning. If these hyper-parameters are not set properly, the models may not converge in the same or close basin. Even based on the pre-trained fine-tuning paradigm, there is still a significant gap between the merged and independent models, especially when the number of models/tasks is large. Therefore, a promising direction for future research is to explore how to ensure the effectiveness of model merging under more relaxed conditions. For example, investigating how to merge multiple models that are trained independently from scratch for different tasks without compromising performance could be valuable.
@@
(2) In-depth Theoretical Analysis for Model Merging. The validity and explanation of existing model merging techniques are largely empirical and lack sufficient theoretical guarantees. As discussed in §2.4, there is currently a limited amount of work on the theoretical aspects or explanations of model merging. The few existing studies mainly focus on merging multiple models trained on the same trajectory or on the same dataset with different fine-tuning settings. There is almost no theoretical research or explanation concerning merging multiple models fine-tuned on different datasets or merging multiple models trained from scratch on different datasets. Therefore, future research should aim for a more comprehensive and in-depth theoretical analysis to enhance the success and reliability of model merging. Furthermore, a deeper understanding of the effectiveness of model merging can, in turn, facilitate the discovery of prior conditions that are more conducive to model merging.
@@
(4) Effective and Efficient Model Merging. Existing high-performance model merging methods often come with significant costs in terms of efficiency and memory. First, most of these methods require all models to be loaded into memory during execution. For instance, merging 72 fine-tuned ViT-B/32 models necessitates more than 200GB of memory [107, 208]. Additionally, heuristics for determining model merging coefficients involve repeated evaluations of the combined model, while learnable methods depend on additional data and training. In the future, it would be beneficial to develop more efficient model merging methods that do not require training, additional data, GPUs, or large amounts of memory.
(5) Merge Heterogeneous Expert Models. Existing methods primarily focus on merging homogeneous models. However, in practice, numerous heterogeneous models excel in various tasks. A limited number of existing methods for merging heterogeneous models involve transforming multiple heterogeneous models into homogeneous ones using knowledge distillation techniques, followed by the merging process [10, 194]. The distillation process relies on the data from the original tasks and involves costly training. Therefore, it is also worth exploring approaches to merge these heterogeneous models without incurring the high costs associated with architectural transformations.
(6) Interdisciplinary Application of Model Merging. As discussed in §3 and §4, model merging has been adeptly applied across various foundation models and machine learning subfields to address different challenges and achieve interesting tasks. The question of how to adapt model merging strategies from one subfield to another presents an exciting avenue for exploration.
6 Conclusions
Model merging is a straightforward and effective technique for model enhancement that combines multiple models to achieve diverse capabilities. In this survey, we first provide a comprehensive overview of the advanced methods and theories currently available in the field of model merging. Next, we discuss the application of model merging techniques across various foundation models (i.e., LLMs, MLLMs) and more than ten subfields of machine learning, highlighting their use in addressing various challenges and difficulties. Finally, we identify ongoing issues within the model merging and propose six research directions that are worthy of further exploration. We believe that model merging technology, as an efficient and modular model empowerment solution, will play a significant role in more practical scenarios in the future.

### v5

@@ 5.3.2. Model Merging for Better Domain Generalization
Domain generalization methods aim to generalize to an unknown target domain using only training data from source domains. For instance, in the context of traffic sign recognition, the training data for a machine learning model tasked with identifying traffic signs in various urban environments comes from multiple cities (i.e., source domains). However, when deployed, the model must recognize traffic signs in new urban environments (i.e., target domains) that it has never encountered before. Existing DG methods can be classified into domain alignment, data augmentation, regularization, and meta-learning frameworks (Arpit et al., 2022). Complementary to these approaches, model merging techniques can be seamlessly integrated to further improve out-of-domain performance without modification (Cha et al., 2021; Arpit et al., 2022; Li et al., 2024b). SWAD (Cha et al., 2021) demonstrates that flatter minima generalize better to unseen domains. Inspired by SWA (Izmailov et al., 2018), SWAD proposes a dense and overfit-aware stochastic weight sampling strategy to identify these flatter minima. More specifically, unlike SWA, it starts from a predefined epoch until the final epoch, and collects a random weight every $K$ epochs for averaging. SWAD collects weights densely, that is, one is collected every step, and the start and end of random weight collection are determined by the performance changes on the validation set. EoA (Arpit et al., 2022) also shows that model averaging can improve out-of-domain stability, and that ensembling multiple moving average models can further enhance performance compared to ensembling models without weight averaging.
5.4. Model Merging in Federated Learning
@@
Federated Learning (FL) is a distributed learning approach that allows multiple clients to collaboratively train a model without sharing data. FL primarily includes two settings: centralized (with a central server) and decentralized (without a central server). Each client updates the model or calculates the gradient based on local data and sends the updated information to the central server (in centralized FL) or other clients (in decentralized FL) for aggregation to update the global model, thus ensuring data privacy protection. In FL, model merging refers to summarizing parameters from various clients during each communication round, thereby forming an updated global model.
@@
Most FL methods adopt a simple coordinate-wise average to aggregate the local models. For example, they calculate local model merging coefficients according to some heuristic rules. FedAvg (McMahan et al., 2017), the most classic FL method, proposes to merge local models on the server weighted by the amount of training data from each client. FedNova (Wang et al., 2020a) normalizes and scales model updates on the client side based on the number of update steps, efficiently aggregating local models to obtain a high-performance global model. FedAtt (Ji et al., 2019) calculates layer-wise attention coefficients based on the similarity of client and server parameters, fusing local models based on these coefficients. FedFisher (Jhunjhunwala et al., 2024b) computes the Fisher information matrix of the parameters in each client to merge the local models. In more challenging FL tasks, the above direct coordinate-wise merging methods may result in suboptimal global model performance. Inspired by the property of permutation invariance of neural networks, PFNM (Yurochkin et al., 2019), OTFusion (Singh and Jaggi, 2020) and FedMA (Wang et al., 2020b) propose to permute neurons of local models before merging them. Similarly, GAMF (Liu et al., 2022a) transforms the merging problem into a multi-graph matching problem based on graph matching and then merges the aligned local models.
5.5. Model Merging in Zero-shot/Few-shot Learning
@@
In practical applications of machine learning models, collecting a large amount of labeled data can be expensive or infeasible in specific scenarios (e.g., medical diagnosis, real-time monitoring). Users often want models to effectively perform new tasks that have not been encountered before, that is, an ability commonly referred to as cross-task generalization (Huang et al., 2024b). Zero-shot (Mishra et al., 2022) and few-shot learning (Ye et al., 2021) can reduce the dependence on large amounts of data and allow the model to better deal with unseen categories or small numbers of samples, improving the cross-task generalization ability of the model. In few-shot learning, a common approach is to fine-tune the model using the limited examples available. However, because of the minimal data, this fine-tuning process is often unstable and yields only modest performance improvements. Recently, some studies have explored merging pre-trained models (from some publicly accessible resources) to enhance cross-task generalization under zero-shot and few-shot conditions (Wu et al., 2023; Huang et al., 2024b; Jang et al., 2023b; Ostapenko et al., 2024; Muqeeth et al., 2024b; Wang et al., 2024c; Tao et al., 2024).
@@
5.5.1. Model Merging for Cross-task Generalization in Zero-shot Learning
@@
Model merging has demonstrated the effectiveness of zero-shot learning across several applications. Some examples of practical applications include cross-lingual transfer (Huang et al., 2024a; Chronopoulou et al., 2023b; Zhao et al., 2024c; Klimaszewski et al., 2024), hybrid style image generation (Biggs et al., 2024; Nair et al., 2024), and multi-modal processing (Chen et al., 2024a). Some works achieve cross-lingual transfer through model merging, such as chat (Huang et al., 2024a), text summarization (Chronopoulou et al., 2023b), or reasoning (Zhao et al., 2024c). A well-performing language-specific LLM needs to be fully trained, and with 7,000 languages in the world, not all of them have enough labeled data to support model fine-tuning. Therefore, cross-lingual knowledge transfer is particularly important. For example, Huang et al. (2024a) build a Chat vector based on fine-tuned LLAMA2-chat and pre-trained LLAMA2 on chat data in the English language, and assembles it with the continuously pre-trained LLAMA2 model on other non-English languages. This allows the new model to chat in non-English languages. Chronopoulou et al. (2023b) develop a zero-shot multilingual summarization framework. It uses a merged model (a supervised summarization model and an unsupervised pre-trained model for a high-resource language, along with an unsupervised pre-trained model for a low-resource language) to perform text summarization tasks in low-resource languages. Similarly, AdaMergeX (Zhao et al., 2024c) demonstrates the effectiveness of model merging for cross-language transfer across three tasks: reasoning, natural language understanding, and natural language generation. In the hybrid-style image generation task, Diffusion Soup (Biggs et al., 2024) and MaxFusion (Nair et al., 2024) show that the zero-shot generation ability can be enhanced by merging multiple diffusion models. In the multi-modality task, DAMC (Chen et al., 2024a) experiments prove that zero-shot multi-modal extension can be achieved by merging multi-modal models, provided they are initialized from the same LLM. For example, by merging a visual LLM and an audio LLM, the combined model can not only perform image or audio tasks independently but also acquire the zero-shot ability to process inputs containing both visual and auditory information simultaneously.
5.5.2. Model Merging for Cross-task Generalization in Few-shot Learning
@@
Parameter-efficient fine-tuning (PEFT), such as LoRA or Adapter, facilitates the creation and sharing of thousands of custom PEFT modules, each trained on different data for various downstream tasks. A natural question is whether combining PEFT modules pre-trained on different upstream tasks can improve the transfer accuracy for unseen downstream tasks with limited samples. Recent work on model merging suggests a positive answer, showing that merged models can enhance generalization in few-shot settings (He et al., 2023a; Asadi et al., 2024; Huang et al., 2024b; Wang et al., 2024c). For example, LoraHub (Huang et al., 2024b) proposes to merge LoRA modules available on HuggingFace to achieve adaptive performance for unseen tasks, where the merging coefficients of different LoRA are searched in a black-box gradient-free manner with few-shot samples. As expected, few-shot LoraHub performs better than few-shot in-context learning and reduces inference costs by eliminating the need for examples as input to LLMs. LoraRetriever (Zhao et al., 2024a) further proposes dynamically retrieving the most relevant LoRAs based on the input and merging them. Similarly, MerA (He et al., 2023a) proposes merging pretrained adapters into a single adapter for few-shot NLP scenarios. In general, well-trained LoRAs or adapters can serve as valuable resources that users can easily share, access, and apply to a variety of downstream tasks. In the real world, upstream and downstream tasks can be entirely disparate, originating from different datasets, domains, or even different parts of the same dataset. Asadi et al. (2024) comprehensively evaluates model merging in the few-shot learning setting. Specifically, this study examines three cases of label, domain, and task drift between upstream and downstream tasks. The results demonstrate that model merging enhances the model’s generalization ability in few-shot learning scenarios across different contexts.
5.6. Model Merging in Adversarial Learning
@@
In the machine learning community, the open-source availability of pre-trained models (e.g., LLaMA (Touvron et al., 2023a)) has accelerated technological advancements. In this context, developers often download unvalidated checkpoints to fine-tune their models or even outsource the training process to third-party platforms (Wu et al., 2022). Consequently, open-source models are also vulnerable to malicious attacks, such as poisoning attacks, where hidden malicious behaviors can be triggered by specific inputs. This raises several intriguing questions: Can model merging lead to attacks, and can it be used to develop defense mechanisms? Additionally, how can intellectual property protection be enhanced in the context of model merging?
5.6.1. Model Merging as an Attack
@@
Parameter-efficient fine-tuning methods (Ding et al., 2023), such as LoRA (Hu et al., 2022), exhibit functional transferability. This means that a LoRA model fine-tuned for a specific task based on a pretrained model can be successfully transferred to another pretrained model (Liu et al., 2024c). In practice, developers often download LoRA models from open-source platforms to address their specific downstream tasks (Huang et al., 2024b). If a poisoned LoRA, which could be seen as a Trojan horse, is inadvertently downloaded and integrated into a model, it may introduce security vulnerabilities. Research by LoRA-as-an-Attack (Liu et al., 2024c) demonstrates that merging a poisoned LoRA—trained on compromised data—with a benign LoRA, trained on clean data, can result in a backdoor injection. This finding also holds when multiple LoRAs are merged. In addition, BadMerging (Zhang et al., 2024b) has developed a two-stage backdoor attack framework specifically for model merging, and through a large number of experiments, it has shown that the success rate of on-task and off-task attacks on merged models exceeds 90%, and existing defense measures cannot defend against BadMerging.
5.6.2. Model Merging as a Defense or Intellectual Property Protection
@@
Contrary to the attacks described in §5.6.1, the transferability of LoRA also offers an opportunity for model merging as a defense strategy. Specifically, if we know that a model may be susceptible to certain attacks, can we train some LoRAs to enhance the model’s defense (i.e., reduce the attacker’s success rate)? For example, Liu et al. (2024c) demonstrate that GPT-3.5 was used to generate a benign dataset containing backdoor triggers. A dedicated defense LoRA was then trained on this benign data and merged into the poisoned pre-trained model. This defensive model merging ultimately led to a reduction in the backdoor effect. Furthermore, research has shown that in the context of full parameter fine-tuning, model merging can serve as a ”free lunch” for model defense. Experiments involving four model architectures and four datasets revealed that merging multiple poisoned models without additional effort mitigated these poisoning attacks, with the accuracy on the benign dataset remaining nearly unaffected. Rebuffi et al. (2023) and Croce et al. (2023) merge a set of $l_{p}$ (for various $p$) robust fine-tuned models to easily control the robustness level of each threat model against $l_{p}$ boundary adversarial attacks. Similarly, the experimental analysis by (Gallego, 2024) indicates that model merging offers an effective defense mechanism against jailbreak attacks.
@@
In another practical scenario, merging unauthorized models may infringe on the intellectual property rights of the model owner. Malicious users might merge several high-quality open-source models (e.g., those authorized for research use only) to create a new model, then claim that this new model was entirely developed and trained from scratch by themselves, subsequently offering model services for commercial gain. In such cases, it becomes particularly crucial for model owners to detect whether others have merged their models. MergeGuard (Cong et al., 2024) performs a preliminary analysis of the effectiveness of two existing defense methods—Quantization Watermarking (Li et al., 2023b) and Instructional Fingerprint (Xu et al., 2024a)—in the context of model merging. The study observed that while the watermarking method cannot be detected in the merged model, the fingerprint method remains detectable.
6. Remaining Challenges and Future Directions
@@
Although §2–§4 and §5 present various advanced model merging methods and applications, challenges remain in the technology and application of existing model merging approaches. Additionally, there are numerous areas that warrant further research in the future.
@@
(1) Closing the Performance Gap Between the Merged and Independent Models. The effectiveness of current model merging techniques heavily relies on the ”pretraining-finetuning” paradigm. Specifically, successful model merging requires that multiple models be fine-tuned based on the same pre-trained model, with careful control over the number of epochs and learning rate during fine-tuning. If these hyperparameters are not set properly, the models may not converge in the same or close basin. Even based on the pre-trained fine-tuning paradigm, there is still a significant gap between the merged and independent models, especially when the number of models/tasks is large. As shown in Figure 6, when merging eight tasks, the independently fine-tuned ViT-B/32 model achieves an accuracy of 90.3, whereas the simplest weight averaging baseline reaches only 66.5. Meanwhile, for Llama-3.1, the best model merging method attains a score of 84.8, which still lags far behind the expert model’s score of 100. Therefore, a promising direction for future research is to explore how to ensure the effectiveness of model merging under more relaxed conditions. For example, investigating how to merge multiple models that are trained independently from scratch for different tasks without compromising performance could be valuable.
@@
(2) In-depth Theoretical Analysis for Model Merging. The validity and explanation of existing model merging techniques are largely empirical and lack sufficient theoretical guarantees. As discussed in §3.1, there is currently a limited amount of work on the theoretical aspects or explanations of model merging. The few existing studies mainly focus on merging multiple models trained on the same trajectory or on the same dataset with different fine-tuning settings. There is almost no theoretical research or explanation concerning merging multiple models fine-tuned on different datasets or merging multiple models trained from scratch on different datasets. Therefore, future research should aim for a more comprehensive and in-depth theoretical analysis to enhance the success and reliability of model merging.
@@
(4) Effective and Efficient Model Merging. Existing high-performance model merging methods often come with high costs in terms of efficiency and memory. First, most of these methods require all models to be loaded into memory during execution. For instance, merging 72 fine-tuned ViT-B/32 models necessitates more than 200GB of memory (Li et al., 2024c; Wortsman et al., 2022a). Additionally, heuristics for determining model merging coefficients involve repeated evaluations of the combined model, while learnable methods depend on additional data and training. In the future, it would be beneficial to develop more efficient model merging methods that do not require training, additional data, GPUs, or large amounts of memory.
(5) Merge Heterogeneous Models. Existing methods primarily focus on merging homogeneous models. However, in practice, numerous heterogeneous models excel in various tasks. A limited number of existing methods for merging heterogeneous models involve transforming multiple heterogeneous models into homogeneous ones using the knowledge distillation technique (Avrahami et al., 2022; Wan et al., 2024b). The distillation process relies on the data from the original tasks and involves costly training. Therefore, it is also worth exploring approaches to merge these heterogeneous models without incurring the high costs associated with architectural transformations.
(6) Develop Merger-Friendly Fine-tuning Strategies. Most existing model merging methods primarily focus on mitigating parameter conflicts and interference at the merging stage. Only a few works consider improving the fine-tuning procedure itself so that the resulting models are more amenable to merging, i.e., they incur smaller performance degradation compared with standard fine-tuning. Designing better fine-tuning strategies that facilitate subsequent merging, without significantly increasing computational cost or imposing strict constraints on developers, is therefore a highly meaningful direction for future work.
(7) Interdisciplinary Application of Model Merging. As discussed in §4 and §5, model merging has been adeptly applied across various foundation models and machine learning subfields to address different challenges and achieve interesting tasks. The question of how to adapt model merging strategies from one subfield to another presents an exciting avenue for exploration. In addition, it is important to further explore additional application scenarios for model merging.
7. Conclusions
Model merging is a straightforward and effective technique for model enhancement that combines multiple models to achieve diverse capabilities. In this survey, we first provide a comprehensive overview of the advanced methods and theories currently available in the field of model merging. Next, we discuss the application of model merging techniques across various foundation models (i.e., LLMs, MLLMs) and more than ten subfields of machine learning, highlighting their use in addressing various challenges and difficulties. Finally, we identify ongoing issues within the model merging and propose seven research directions that are worthy of further exploration. We believe that model merging technology, as an efficient and modular model empowerment solution, will play a significant role in more practical scenarios in the future.